two ressistance Xand Y arr in left and right meter bridge the balance point is 40 cm from left of 10 resistivety each are connected series with X AND Y seprately.the balance point is 45 cm . the value of X and Y
 

Grade:9

1 Answers

Neeti
8 years ago
X = 8 and Y = 12. 
in first case, X/40=Y/60
In second case,  (X+10)/45= (Y+10)/55
solving these two by simultaneous equations, you get the answer :) 
please like and approve :)